greg has two soulutions avaiable in a lAB, one with 6% alcohol and another with 14% alcohol. how much of each should he mix together to obtain 10 gallons of a solution hat contains 12% alcohol.
Answer by mananth(16949)

one with 6% alcohol : let this be x gallons
another with 14% alcohol. this will be 10-x gallons
mix 10 gallons 12% alcohol.
...
6x+14(10-x)=10*12
6x+140-14x=120
-8x=-20
x=2.5 gallons 6% alcohol
7.5 gallons 14% alcohol
...
m.ananth@hotmail.ca